This is my first time on this website, so I'm sorry if I chose the wrong subject, but here's my question. How would you solve the answer to Y/3-4=12 ?
So it is the fraction Y over 3 subtract 4 equals 12. I need the steps to solve it not just the answer, so when I have future problems like this I'll know how to do it. Thank you for the help! And also in the problem you must write y as a function of x and evaluate when x is -2, 0, and 2. Thanks again!

Answer by checkley77(12844)

Y/3-4=12 ADD 4.
Y/3=12+4 COMBINE INTEGERS.
Y/3=16 CROSS MULTIPLY.
Y=16*3
Y=48 ANS.
PROOF:
48/3-4=12
16-4=12
12=12
------------------
EXAMPLE: Y=4X
WHEN X=-2,0,2
Y=4*2=8
Y=4*0=0
Y=4-2=-8